Using the information given, put each line in Slope-intercept form.
Slope 2: 7-intercept; 5
Slope 1/2; x-intercept; 7

Answer by venugopalramana(3286)

Using the information given, put each line in Slope-intercept form.
Slope 2: 7-intercept; 5...HOPE YOU MEAN Y INTERCEPT IS 5..IF..SO...
Y=MX+C...IS THE STD.EQN.WHERE M IS SLOPE AND C IS Y INTERCEPT..SO EQN. IS
Y=2X+5
Slope 1/2; x-intercept; 7
Y=MX+C....Y=X/2+C....X INTERCEPT IS 7 MEANS WHEN Y=0..X=7
0=7/2+C.....C=-7/2...HENCE EQN.IS ..
Y=X/2-7/2....OR.....2Y=X-7